Bill is three years older than Ted. In five years, the sum of their ages will be 47. How old are they now
Answer by JulietG(1812)

B = T + 3 [Bill is three years older than Ted.]
(B+5) + (T+5) = 47 [In five years, the sum of their ages will be 47.]
B + T = 37 (5 years and 5 years from the previous equation)
Substitute the value of B in the first equation into the third.
(T+3)+ T = 37
Add
2T + 3 = 37
Subtract 3 from each side
2T = 34
Divide each side by 2
T = 17
If Ted is 17, then Bill is 20.
In five years, Ted will be 22 and Bill will be 25.
22 + 25 = 47